# Surface-Aware Distilled 3D Semantic Features

### v1

@@ Abstract
Many 3D tasks such as pose alignment, animation, motion transfer, and 3D reconstruction rely on establishing correspondences between 3D shapes. This challenge has recently been approached by matching of semantic features from pre-trained vision models. However, despite their power, these features struggle to differentiate instances of the same semantic class such as “left hand” versus “right hand” which leads to substantial mapping errors. To solve this, we learn a surface-aware embedding space that is robust to these ambiguities. Importantly, our approach is self-supervised and requires only a small number of unpaired training meshes to infer features for new 3D shapes at test time. We achieve this by introducing a contrastive loss that preserves the semantic content of the features distilled from foundational models while disambiguating features located far apart on the shape’s surface. We observe superior performance in correspondence matching benchmarks and enable downstream applications including in-part segmentation, pose alignment, and motion transfer.
Figure 1: We introduce a surface-aware feature embedding space separating instances of the same semantic class (e.g. right vs. left instances for “hand/paw”). Our method implicitly aligns semantically-related regions (shown with same colors) across diverse 3D shapes in a self-supervised manner without access to any correspondence labels. See Sec. 6.1 for figure details. The project site is available at: https://lukas.uzolas.com/SurfaceAware3DFeaturesSite.
1 Introduction
@@
Establishing accurate correspondence between 3D shapes is important in many applications, including 3D registration, pose alignment, motion transfer, as well as static and dynamic 3D reconstruction. Historically, geometric descriptors have been used to determine matches between 3D shapes under isometric deformations, but they struggle with non-isometric deformations [72, 76, 5]. In contrast, neural features, stemming from pre-trained 2D vision models, have recently achieved great success in identifying correspondences between vastly different shapes [75, 51, 92, 81], such as mapping from cats to lions.
@@
Despite this inter-class robustness, neural features often struggle to disambiguate between instances of the same class like “left hand” and “right hand” (see  Fig. 6). Such mismatches can lead to substantial errors in downstream applications (see Sec. 6.2). Recent research has demonstrated that these features contain global pose information and that disambiguation is possible in a 2D scenario [91]. However, achieving the same effect on distilled 3D features is not trivial, especially in a low-data regime, which is prevalent in 3D, where data acquisition and labeling is difficult.
@@
Our work improves 3D neural features distilled from pre-trained 2D vision models by embedding them into a space disambiguating intraclass instances. We achieve this without large annotated datasets using a self-supervised learning scheme guided by in-shape geodesic distances without the need for shape pairs. Training with a limited number of 3D meshes, our method produces surface-aware features for diverse new shapes. In quantitative and qualitative comparisons to prior work, we demonstrate superior suitability of these features to serve as robust descriptors for matching and as building blocks for solving other tasks.
@@
In summary, we make the following contributions: 1. We introduce a novel contrastive loss for self-supervised distillation of 3D features. 2. We demonstrate the effectiveness of our surface-aware features in multiple experiments and correspondence matching benchmarks. 3. We showcase two additional downstream applications that motivate further future use cases.
2 Related Works
@@
Our method utilizes contrastive learning to improve semantic features extracted from foundational models for robust 3D shape correspondence matching. In this section, we discuss prior work in these three areas.
2.1 Shape correspondences
@@
Classical shape registration methods directly minimize global [7] or local [9] inter-shape distances making them susceptible to local minima [85]. This motivates the design of more informative local geometric descriptors [72, 76, 5]. These can alternatively be learned [16, 36] from voxelixed patches [88, 33, 4] or from point clouds [28, 87, 20, 19, 21]. The learning can be supervised by labels [16] or achieved without them [28, 35, 44, 89]. Our method falls into the latter category, as our contrastive loss motivates our encoder to separate instances by approximating geodesic distances [83] without training data labels. This is conceptually similar to previous methods for near-isometric shape deformations [54, 38, 67]. However, we distinctly do not measure geodesic distortions between shape pairs, and therefore we do not limit our method to isometric deformations, and we do not compute any geodesics during inference. Instead, we only use the geodesics to disambiguate information already available in the image-based features, which is critical for our results.
The correspondences can be recovered from descriptors by a matching [31], directly regressed [50, 79] or established on parametric templates [35, 22]. Instead, we focus on the descriptors themselves, and we show several different application scenarios in Sec. 6.
@@
Functional Maps (FMs) [58] allow for matching on a surface. FMs are real-valued surface functions in the space of Laplace-Bertrami eigenfunctions, supporting linear transformations between shapes. Constrained to match surface descriptors for each shape [72, 76, 5] they allow extracting point-wise correspondences [58, 63]. These functions can also be learned [46] often with little or no supervision [65, 23, 32, 38, 73]. Extrinsic alignment can support nonisotropic deformations [26, 27]. In this work, we focus on improving features for direct point-to-point matching in the spatial domain, but previous work demonstrated promising results when combining similar image-based features with FMs [55, 24].
2.2 Image-based features for 3D shapes
Image-based features emerge in large visual models for 2D image tasks. Self-supervised features from Vision Transformers such as DINO-ViT [12] and DINOv2 [57] locally encode semantic information useful for segmentation [12] or image-to-image correspondence matching [2]. SD-DINO [90] adds complementary features from the diffusion-based image synthesis model Stable Diffusion [64]. Lifting these features to 3D has enabled the self-supervised construction of canonical surface maps [70], transfer of appearance between 3D shapes [30], 3D animation [78], keypoint detection [81] or matching of surface correspondences [24, 55, 14]. However, despite their semantic versatility, disambiguating between intraclass instances, such as left and right hands, remains challenging but possible, as shown in a recent 2D image study [91] This motivates our 3D shape descriptors for resolving instance ambiguity. We evaluate the Diff3F features [24] as a basis for our autoencoded embedding. We then use geodesics as a guiding signal, which is conceptually similar to the concurrent work [14]. However, we demonstrate that our distinct design choices lead to better performance in Sec. 5.3. Moreover, our method is purely intrinsic and does not require aligned mesh pairs for training.
Beyond vision-only models, multimodal large language models have recently been effective in image and 3D shape analysis including keypoint labeling [34] and shape co-segmentation [1]. In our work, we focus on vision-only models because of their simplicity, but we consider a model combination a promising research direction.
2.3 Contrastive Learning
Contrastive learning embeds similar samples close to each other while keeping dissimilar samples apart. This can be achieved directly by minimizing and maximizing embedding distances for positive and negative pair samples, respectively [15, 37, 80, 66] or indirectly, such as by optimizing performance in an auto-regressive task [56]. Training pairs can be obtained by data augmentation [13], from memory banks [82, 40], or by clustering [10, 11]. Learning with cross-domain labels yields joint embeddings, as demonstrated by CLIP [62] for text and images. We design our contrastive loss to disambiguate intraclass instances guided by a geodesic metric.
@@ 2.3 Contrastive Learning
Figure 2: Overview of our method. We feed images of a 3D shape rendered from multiple viewpoints to a pre-trained 2D vision model and extract features that are then projected back onto surface points $\mathbf{p}_{i}$ and aggregated into per-point features $\mathbf{f}_{i}$ (Sec. 3). Next, we point-wise embed the base features $\mathbf{f}_{i}$ into our surface-aware features $\mathbf{s}_{i}$ residing in a lower-dimensional space learned using our contrastive loss preserving geodesic distances $d_{i,j}$ and a reconstruction loss matching decoded features $\mathbf{\bar{f}}_{i}$ to $\mathbf{f}_{i}$ (Sec. 4). The surface-aware features $\mathbf{s}_{i}$ serve as robust descriptors for correspondence matching (Sec. 5) and base blocks for other tasks (Sec. 6).
3 Preliminaries
@@
We built upon methods that aggregate features from pre-trained 2D vision models on 3D meshes [24, 81, 55, 14]. In this section, we give a brief overview on these methods.
3.1 Reprojection of 2D Features
@@
We represent a 3D shape as a triangular mesh with a tuple of $N$ vertices and $M$ triangular faces, that is, :=($\mathcal{M}{}:=(\{\mathbf{p}_{n}\in\mathbb{R}^{3}|n=1,...,N\},\{\mathbf{t}_{m}\in\mathbb{N}^{3}|m=1,...,M\})$. The rendering function $R_{rgb}:(\mathcal{M},\mathcal{C})\rightarrow\mathbf{I}_{rgb}$ projects $\mathcal{M}{}$ into a camera $\mathcal{C}$ and outputs an image $\mathbf{I}_{rgb}\in\mathbb{R}^{H\times W\times 3}$, with height $H$ and width $W$. Optionally, texturing is possible in $R_{rgb}(.)$ or as a ControlNet [93] post-processing. The image is then passed to a pre-trained vision model [12, 57, 64, 90] to obtain dense semantic feature maps $\mathbf{F}\in\mathbb{R}^{h\times w\times f}$ with $h,w,f$ as two spatial and one feature dimension. Finally, per-vertex features $\mathbf{f}_{n}\in\mathbb{R}^{f}$ are obtained by projective texture mapping of $\mathbf{F}$ onto $\mathcal{M}$. To cover the whole surface, features are aggregated across multiple cameras, resulting in a set of features :=$\mathcal{F}_{\mathcal{M}}:=\{\mathbf{f}_{n}\in\mathbb{R}^{3}|n=1,...,N\}$. Throughout this work, we refer to $\mathcal{F}_{\mathcal{M}}$ as the base features on which our method is built. The exact choice of $\mathcal{F}_{\mathcal{M}}$ is orthogonal to our contribution but must encode semantic information. To this extent, we use Diff3F [24] in this work.
@@
Features $\mathcal{F}_{\mathcal{M}}$ have been shown to encode strong semantic information useful for correspondence matching [75, 24]. In the simplest case, the feature $\mathbf{f}_{n}\in\mathcal{F}_{\mathcal{T}}$ of a target mesh $\mathcal{T}$ that best matches the feature $\mathbf{f}_{m}\in\mathcal{F}_{\mathcal{S}}$ of a source mesh $\mathcal{S}$ is determined by maximizing the cosine similarity $\phi:\mathbb{R}^{f}\times\mathbb{R}^{f}\rightarrow\mathbb{R}$:
$\phi(\mathbf{f}_{i},\mathbf{f}_{j})=\frac{\mathbf{f}_{i}^{T}\mathbf{f}_{j}}{\lVert\mathbf{f}_{i}\lVert_{2}\lVert\mathbf{f}_{j}\lVert_{2}},$ (1)
such that $\tau(\mathbf{p}_{m})=\arg\max_{\mathbf{p}_{n}}\phi(\mathbf{p}_{n}\to\mathbf{f}_{n},\mathbf{p}_{n}\to\mathbf{f}_{m})$ is the best matching point. However, the features $\mathcal{F}_{\mathcal{M}}$ do not differentiate between semantic instances well (see  Fig. 5) which we address by learning robust surface-aware features $\mathcal{S}_{\mathcal{M}}$.
4 Method

### v2

@@ Abstract.
Many 3D tasks such as pose alignment, animation, motion transfer, and 3D reconstruction rely on establishing correspondences between 3D shapes. This challenge has recently been approached by pairwise matching of semantic features from pre-trained vision models. However, despite their power, these features struggle to differentiate instances of the same semantic class such as “left hand” versus “right hand” which leads to substantial mapping errors. To solve this, we learn a surface-aware embedding space that is robust to these ambiguities while facilitating shared mapping for an entire family of 3D shapes. Importantly, our approach is self-supervised and requires only a small number of unpaired training meshes to infer features for new possibly imperfect 3D shapes at test time. We achieve this by introducing a contrastive loss that preserves the semantic content of the features distilled from foundational models while disambiguating features located far apart on the shape’s surface. We observe superior performance in correspondence matching benchmarks and enable downstream applications including 2D-to-3D and 3D-to-3D texture transfer, in-part segmentation, pose alignment, and motion transfer in low-data regimes. Unlike previous pairwise approaches, our solution constructs a joint embedding space, where both seen and unseen 3D shapes are implicitly aligned without further optimization. The code is available at https://graphics.tudelft.nl/SurfaceAware3DFeatures.
Semantic Features, Contrastive Learning, Motion Transfer, Reposing, Shape Correspondences
$c$s: Computing methodologies Shape analysisccs: Computing methodologies Dimensionality reduction and manifold learningccs: Computing methodologies Motion processing
Figure 1. We introduce a surface-aware feature embedding space based on 2D pre-trained foundational models. In contrast to related works, our embedding space separates instances of the same semantic class (e.g. right vs. left instances for “hand/paw”) which facilitates many downstream applications: a) Texturing of even incomplete 3D shapes based on a 2D image (the input image was produced by ChatGPT), b) 3D correspondences between non-isometric shapes, c) Re-posing of meshes based on a single source and target pair, d) Pose alignment of 3D meshes with dense and sparse point correspondences. :
1. Introduction
@@
Mapping 3D shapes into a shared space guaranteeing mutual correspondences is important for many applications, including 3D registration, pose alignment, motion transfer, as well as static and dynamic 3D reconstruction. Historically, geometric descriptors have been used to determine matches between pairs of 3D shapes under isometric deformations, but they struggle with non-isometric deformations (Sun et al., 2009; Tombari et al., 2010; Aubry et al., 2011). In contrast, neural features, stemming from pre-trained 2D vision models, have recently achieved great success in identifying correspondences between pairs of vastly different shapes (Tang et al., 2023; Luo et al., 2023; Zhang et al., 2024b; Wimmer et al., 2024), such as mapping from cats to lions. In this paper, we make another important step by moving from one-to-one pairwise shape correspondence matching to a joint embedding space establishing many-to-many shape correspondences.
@@
Despite their inter-class robustness, neural features often struggle to disambiguate between instances of the same class like “left hand” and “right hand” (see  Fig. 3). Such mismatches can lead to substantial errors in downstream applications (see Sec. 6). Recent research has demonstrated that these features contain global pose information and that disambiguation is possible in a 2D scenario (Zhang et al., 2024a). However, achieving the same effect on distilled 3D features is not trivial, especially in a low-data regime, which is prevalent in 3D, where data acquisition and labeling is difficult.
@@
Our work improves 3D neural features distilled from pre-trained 2D vision models by embedding them into a space disambiguating intraclass instances. We achieve this without large annotated datasets using a self-supervised learning scheme guided by in-shape geodesic distances without the need for shape pairs. Training with a limited number of 3D meshes, our method produces space of surface-aware features establishing multi-faceted correspondences for diverse new shapes without any further fine-tuning. In quantitative and qualitative comparisons to prior work, we demonstrate superior suitability of these features to serve as robust descriptors for matching and as building blocks for solving other tasks. Since geodesic distances are not used during inference, our method has only a minimal overhead from its shallow neural encoder and its point-wise nature makes it robust to varying mesh complexity or shape incompleteness. Finally, the encoder preserves compatibility with per-pixel image features, and hence, also naturally establishes robust 2D-to-3D mappings.
@@
In summary, we make the following contributions: 1. We introduce a novel contrastive loss for self-supervised distillation of 3D features. 2. We quantitatively demonstrate the effectiveness of our surface-aware features in pose transfer, correspondence matching and skinning weight regression. 3. We showcase versatility of our approach in additional downstream applications including pose alignment, instance-based part segmentation and 2D-to-3D or 3D-to-3D texture transfer. 4. We show the versatility of our features when matching many-to-many shapes of not only humanoids and animals but also other classes.
2. Related Works
@@
Our method utilizes contrastive learning to embed semantic features from foundational models to a space enabling robust n-to-n 3D shape matching. In this section, we discuss prior work in these three areas.
2.1. Image-based features for 3D shapes
Image-based features emerge in large visual models for 2D image tasks. Self-supervised features from Vision Transformers, such as DINO-ViT (Caron et al., 2021) and DINOv2 (Oquab et al., 2024), locally encode semantic information useful for segmentation (Caron et al., 2021) or image-to-image correspondence matching (Amir et al., 2021). SD-DINO (Zhang et al., 2023a) adds complementary features from the diffusion-based image synthesis model Stable Diffusion (Rombach et al., 2022). Lifting these features to 3D has enabled the self-supervised construction of canonical surface maps (Shtedritski et al., 2024), transfer of appearance between 3D shapes (Fischer et al., 2024), 3D animation (Uzolas et al., 2024), keypoint detection (Wimmer et al., 2024) or matching of surface correspondences (Dutt et al., 2024; Morreale et al., 2024; Chen et al., 2025). However, despite their semantic versatility, disambiguating between intraclass instances, such as left and right hands, remains challenging but possible, as shown in a recent 2D image study (Zhang et al., 2024a). This motivates our 3D shape descriptors for resolving in- stance ambiguity.
Prior work tackled this ambiguity by mapping shapes to a spherical template (Mariotti et al., 2024), which is difficult for complex shapes including humans. Alternatively, Liu et al. (2025) recovered non-isometric correspondences from a 2D semantic flow learned from vision features. Instead, we adapt Diff3F features (Dutt et al., 2024) in 3D space and resolve the ambiguity through contrastive learning enforcing geodesic distances.
Geodesic distances have previously supported point cloud analysis (He et al., 2019) and more recently, NIE (Jiang et al., 2023) and the concurrent work DV-Matcher (Chen et al., 2025) similarly utilize geodesic distances for feature disambiguation. Yet, the previously mentioned methods rely on aligned mesh pairs or a learned alignment, while our method learns purely from intrinsic properties of individual shapes. This eases adaptation to less common classes beyond humanoids and animals (Fig. 8) and cross-class mappings (Fig. 9).
Beyond vision-only models, multimodal large language models have recently been effective in image and 3D shape analysis including keypoint labeling (Gong et al., 2024) and shape co-segmentation (Abdelreheem et al., 2023). In our work, we focus on vision-only models because of their simplicity, but we consider a model combination a promising research direction.
2.2. Contrastive Learning
Contrastive learning embeds similar samples close to each other while keeping dissimilar samples apart. This can be achieved directly by minimizing and maximizing embedding distances for positive and negative pair samples, respectively (Chopra et al., 2005; Hadsell et al., 2006; Weinberger and Saul, 2009; Schroff et al., 2015) or indirectly, such as by optimizing performance in an auto-regressive task (Oord et al., 2018). Training pairs can be obtained by data augmentation (Chen et al., 2020), from memory banks (Wu et al., 2018; He et al., 2020), or by clustering (Caron et al., 2018, 2020). Learning with cross-domain labels yields joint embeddings, as demonstrated by CLIP (Radford et al., 2021) for text and images. Contrastive learning was applied to learn end-to-end pose transfer from multiple unregistered meshes of the same identity in different poses (Sun et al., 2023a). We design our contrastive loss to disambiguate intra-class instances guided by a geodesic metric, while learning from intrinsic properties of individual meshes rather than same-identity shape pairs.
2.3. Shape correspondences
@@
Classical shape registration methods directly minimize global (Besl and McKay, 1992) or local (Brown and Rusinkiewicz, 2007) inter-shape distances making them susceptible to local minima (Yang et al., 2015). This motivates the design of more informative local geometric descriptors (Sun et al., 2009; Tombari et al., 2010; Aubry et al., 2011). These can alternatively be learned (Corman et al., 2014; Guo et al., 2015) from voxelized patches (Zeng et al., 2017; Gojcic et al., 2019; Attaiki et al., 2023) or from point clouds (Elbaz et al., 2017; Yew and Lee, 2018; Deng et al., 2018b, a, 2023). The learning can be supervised by labels (Corman et al., 2014) or achieved without them (Elbaz et al., 2017; Groueix et al., 2018; Lang et al., 2021; Zeng et al., 2021). Our method falls into the latter category, as our contrastive loss motivates our encoder to separate instances by approximating geodesic distances (Xia et al., 2021) without training data labels. This is conceptually similar to previous methods for near-isometric shape deformations (Mémoli and Sapiro, 2005; Halimi et al., 2019; Shamai and Kimmel, 2017). However, we distinctly do not measure geodesic distortions between shape pairs, and therefore we do not limit our method to isometric deformations, and we do not compute any geodesics during inference. Instead, we only use the geodesics to disambiguate information already available in the image-based features, which is critical for our results.
The correspondences can be recovered from descriptors by a matching (Fischler and Bolles, 1981), directly regressed (Lu et al., 2019; Wang and Solomon, 2019) or established on parametric templates (Groueix et al., 2018; Deprelle et al., 2019). Here, we focus on the descriptors themselves, and we show several different application scenarios in Sec. 6.
@@
Functional Maps (FMs) (Ovsjanikov et al., 2012) allow for matching on a surface. FMs are real-valued surface functions in the space of Laplace-Bertrami eigenfunctions, supporting linear transformations between shapes. Constrained to match surface descriptors for each shape (Sun et al., 2009; Tombari et al., 2010; Aubry et al., 2011) they allow extracting point-wise correspondences (Ovsjanikov et al., 2012; Rodolà et al., 2015). These functions can also be learned (Litany et al., 2017) often with little or no supervision (Roufosse et al., 2019; Donati et al., 2020; Ginzburg and Raviv, 2020; Halimi et al., 2019; Sun et al., 2023b). Extrinsic alignment can support nonisotropic deformations (Eisenberger et al., 2020a, b). In this work, we focus on improving features for direct point-to-point matching in the spatial domain, but we later demonstrate a combination of our features with FM.
@@
Figure 2. Overview of our method. We feed images of a 3D shape rendered from multiple viewpoints to a pre-trained 2D vision model and extract features that are then projected back onto surface points $\mathbf{p}_{i}$ and aggregated into per-point features $\mathbf{f}_{i}$ (Sec. 3). Next, we pointwise embed the base features $\mathbf{f}_{i}$ into our surface-aware features $\mathbf{s}_{i}$ residing in a lower-dimensional space learned using our contrastive loss preserving geodesic distances $d_{i,j}$ and a reconstruction loss matching decoded features $\mathbf{\bar{f}}_{i}$ to $\mathbf{f}_{i}$ (Sec. 4). The surface-aware features $\mathbf{s}_{i}$ serve as robust descriptors for correspondence matching (Sec. 5) and base blocks for many down-stream applications (Sec. 6).
3. Preliminaries
@@
We build upon methods that aggregate features from pre-trained 2D vision models on 3D meshes (Dutt et al., 2024; Wimmer et al., 2024; Morreale et al., 2024; Chen et al., 2025). In this section, we give a brief overview on these methods.
3.1. Reprojection of 2D Features
@@
We represent a 3D shape as a triangular mesh with a tuple of $N$ vertices and $M$ triangular faces, that is, $\mathcal{M}{}:=(\{\mathbf{p}_{n}\in\mathbb{R}^{3}|n=1,...,N\},\{\mathbf{t}_{m}\in\mathbb{N}^{3}|m=1,...,M\})$. The rendering function $R_{rgb}:(\mathcal{M},\mathcal{C})\rightarrow\mathbf{I}_{rgb}$ projects $\mathcal{M}{}$ into a camera $\mathcal{C}$ and outputs an image $\mathbf{I}_{rgb}\in\mathbb{R}^{H\times W\times 3}$, with height $H$ and width $W$. Optionally, texturing is possible in $R_{rgb}(.)$ or as a ControlNet (Zhang et al., 2023b) post-processing. The image is then passed to a pre-trained vision model (Caron et al., 2021; Oquab et al., 2024; Rombach et al., 2022; Zhang et al., 2023a) to obtain dense semantic feature maps $\mathbf{F}\in\mathbb{R}^{h\times w\times f}$ with $h,w,f$ as two spatial and one feature dimension. Finally, per-vertex features $\mathbf{f}_{n}\in\mathbb{R}^{f}$ are obtained by projective texture mapping of $\mathbf{F}$ onto $\mathcal{M}$. To cover the whole surface, features are aggregated across multiple cameras, resulting in a set of features $\mathcal{F}_{\mathcal{M}}:=\{\mathbf{f}_{n}\in\mathbb{R}^{f}|n=1,...,N\}$. Throughout this work, we refer to $\mathcal{F}_{\mathcal{M}}$ as the base features on which our method is built. The exact choice of $\mathcal{F}_{\mathcal{M}}$ is orthogonal to our contribution but must encode semantic information. To this extent, we use Diff3F (Dutt et al., 2024) in this work.
@@
Features $\mathcal{F}_{\mathcal{M}}$ have been shown to encode strong semantic information useful for correspondence matching (Tang et al., 2023; Dutt et al., 2024). In the simplest case, the feature $\mathbf{f}_{n}\in\mathcal{F}_{\mathcal{T}}$ of a target mesh $\mathcal{T}$ that best matches the feature $\mathbf{f}_{m}\in\mathcal{F}_{\mathcal{S}}$ of a source mesh $\mathcal{S}$ is determined by maximizing the cosine similarity $\phi:\mathbb{R}^{f}\times\mathbb{R}^{f}\rightarrow\mathbb{R}$:
(1) $\phi(\mathbf{f}_{i},\mathbf{f}_{j})=\frac{\mathbf{f}_{i}^{T}\mathbf{f}_{j}}{\lVert\mathbf{f}_{i}\lVert_{2}\lVert\mathbf{f}_{j}\lVert_{2}},$
such that $\tau(\mathbf{p}_{m})=\arg\max_{\mathbf{p}_{n}}\phi(\mathbf{p}_{n}\to\mathbf{f}_{n},\mathbf{p}_{m}\to\mathbf{f}_{m})$ is the best matching point. However, the features $\mathcal{F}_{\mathcal{M}}$ do not differentiate between semantic instances well (see  Fig. 8) which we address by learning robust surface-aware features $\mathcal{S}_{\mathcal{M}}$.
4. Method